# II. Carbon-enhanced metal-poor starsthanks: All M3DIS 3D RHD CEMP models are available under https://nlte.mpia.de

### v1

@@ Abstract
Aims. Understanding the origin and evolution of carbon-enhanced metal-poor (CEMP) stars is key to tracing the early chemical enrichment of the Galaxy. In this work, we investigate how physically realistic 3D radiation-hydrodynamics (RHD), carbon-enhanced model atmospheres affect the inferred carbon abundances in CEMP stars, and assess the implications for their classification and for the Galactic chemical evolution (GCE). We pay particular attention to the systematic biases introduced by traditional 1D hydrostatic equilibrium (HE) models.
Methods. We use the M3DIS code to compute 3D RHD model atmospheres for main-sequence and sub-giant stars spanning a wide range of metallicities and carbon enhancements. Synthetic spectra of the CH G-band are calculated using full 3D radiative transfer, and compared to spectra from classical 1D HE MARCS models. We derive abundance corrections and apply them to a large literature sample of metal-poor stars from the SAGA database to quantify systematic effects on the carbon abundance distribution and CEMP classification.
Results. Our new 3D CEMP models predict significantly cooler upper atmospheric layers than present in 1D HE models, resulting in stronger CH absorption and lower inferred carbon abundances by up to 0.9-0.9 dex at the lowest metallicities. Carbon enhancement in the atmosphere itself increases molecular opacities and leads to radiative re-heating, which partly offsets the adiabatic cooling in 3D models and reduces the magnitude of $\rm 3D-1D$ abundance corrections. Applying these corrections lowers the CEMP fraction by up to $20\%$ below $\rm[Fe/H]=-3$, and furthermore alters the relative contribution of CEMP sub-classes. In particular, the fraction of CEMP-no stars increases while the number of stars classified as CEMP-r/s decreases, owing to the downward revision of absolute carbon abundances. These changes bring the Galactic carbon abundance distribution into better agreement with GCE models that assume a contribution from faint supernovae of $20\%$. Physically realistic model atmospheres are thus essential for reliable reconstruction of the early chemical enrichment history of the Galaxy.
@@ 1 Introduction
Carbon-Enhanced Metal-Poor (CEMP) stars are of major interest in modern astrophysics. Chemical abundances measured in spectra of these stars are relevant for understanding the properties of type II supernovae events (SN II) in the early Galaxy, nucleosynthesis, formation and evolution of binary stars, the stellar initial mass function, the metallicity distribution function of the Galactic halo, and the production sites of neutron-capture elements (Beers & Christlieb 2005; Aoki et al. 2006; Hansen et al. 2018).
@@
CEMP stars are defined based on the amount of carbon (C) compared to iron (Fe) relative to the solar composition present in their atmospheres. Originally, the limit for CEMP classification was placed at $\rm[C/Fe]>1.0$ (Beers & Christlieb 2005), which was later slightly adjusted to the presently applied criterion of $\rm[C/Fe]>0.7$ (Aoki et al. 2007). It is furthermore common to distinguish CEMP stars by categories based on the presence of elements synthesized in the slow- or rapid neutron-capture processes (s- and r-process, Hirschi et al. 2006). The categories include stars with enhancements in either of the two channels exclusively (CEMP-s and CEMP-r, respectively), both (CEMP-r/s), or none (CEMP-no), and each category is thought to probe a different formation environment (e.g. Masseron et al. 2010). CEMP-s stars are viewed as products of enrichment by a post-AGB companion in a binary system (McClure & Woodsworth 1990; Lucatello et al. 2005; Starkenburg et al. 2013; Hansen et al. 2016a), although some CEMP-s stars appear to be single, which might indicate that they formed from the ejecta of a massive, fast rotating source star with a late mixing event (Choplin 2017). In contrast, CEMP-r stars show strong enhancement of r-process elements, such as Eu (Sneden et al. 2003), and are thought to form from matter that became chemically polluted by either a core-collapse supernova (electron-capture or magneto-rotating), neutron star – neutron star, or neutron star – black hole mergers (Frebel & Norris 2015; Côté et al. 2017; Hansen et al. 2018; Kobayashi et al. 2020).
@@
The formation history of CEMP-r/s and CEMP-no stars is less clear. Regarding CEMP-r/s, there is evidence for an r-process contribution (Aoki et al. 2006) from a SN II. Alternative explanations include s-process enrichment by a thermally pulsating AGB companion (Masseron et al. 2006) and neutrino-driven winds from the accretion-induced collapse of an O-Ne-Mg white dwarf (Cohen et al. 2003; Jonsell et al. 2006; Kobayashi et al. 2020). CEMP-no stars are among the most metal-poor stars observed (Christlieb et al. 2002; Hansen et al. 2016b) and they show no significant enrichment in either r- or s-process (Spite et al. 2013). This can be interpreted in the framework of different scenarios, including nucleosynthesis in faint low-energy SNe (Umeda et al. 2005; Tominaga et al. 2007; Nomoto et al. 2013; Kobayashi et al. 2020), hypernovae (Heger & Woosley 2002; Limongi et al. 2003; Nomoto et al. 2013), nucleosynthesis in C-rich winds of rotating massive stars (Hirschi et al. 2006; Meynet et al. 2006), and early-AGB stars (Masseron et al. 2010, and references therein).
@@
The analysis of C abundances in the atmospheres of CEMP stars primarily rests upon spectroscopic diagnostic of the so-called G-band (e.g. Placco et al. 2016; Placco & Beers 2016; François et al. 2018; Norris & Yong 2019). The G-band represents the fundamental A-X transition of the CH molecule (Masseron et al. 2014; Popa et al. 2023) and is therefore highly sensitive to the abundance of C and to the temperature gradient in stellar atmospheres. The G-band has been extensively studied using 1D LTE models (e.g. Lai et al. 2007; Roederer et al. 2014; Jacobson et al. 2015; Lee et al. 2013, see also Suda et al. 2008, 2017 and references therein), but recent analyses also explored the formation of the G-band in 3D radiation-hydrodynamics (RHD) model atmospheres of metal-poor dwarfs and red giants (e.g. Collet et al. 2007; Bonifacio et al. 2013; Gallagher et al. 2016, 2017; Collet et al. 2018).
@@
In this paper, we investigate the effect of carbon enhancement on the structure of stellar atmospheres in realistic 3D RHD simulations introduced in our previous work (Eitner et al. 2024). We further study and quantify the impact of 3D convection on diagnostics of C abundances from stellar spectra. We then analyse the results with respect to CEMP statistics in the Galaxy and discuss implications for the formation and evolution of CEMP stars.
@@
The paper is organised as follows. In Sec. 2, we briefly introduce the stellar observations used in this work, followed by a description of the M3DIS code and model parameters. In Sec. 3, we present our new 3D RHD model atmosphere calculations, including a comparison with 1D hydrostatic equilibrium (HE) MARCS models, the spectrum synthesis of the CH G-band between 42974297\,Å and 43034303\,Å, and the analysis of CEMP abundances in 3D, along with a discussion of their implications for Galactic CEMP populations. Finally, in Sec. 4, we summarise our findings and discuss their implications for the study of CEMP stars and Galactic chemical evolution (GCE).
@@ 2 Methods
Table 1: Summary of M3DIS radiation-hydrodynamics code features.
@@
Radiative Transfer Short-characteristics; conservative on staggered mesh; 8 frequency bins; 4 $\rm\phi$, 2 $\rm\mu$ angles, plus disk-center rays
@@ 2.1 Stellar observations
In this paper we rely on stellar C abundances provided in the Stellar Abundances for Galactic Archaeology database [^1] http://sagadatabase.jp/ (SAGA, Suda et al. 2008, 2011, 2017). SAGA is a comprehensive source of abundances and stellar parameters for stars in the Milky Way and its satellites, with a particular focus on metal-poor stars. The data is based on published, peer-reviewed sources and it has a particular focus on homogeneity of measurements, such that abundances are derived from 1D LTE synthetic spectra (Suda et al. 2008).
The focus of this paper is on main-sequence (MS) and sub-giant stars. Hence, we do not include stars with a surface gravity $\rm log(g)<2.7$. For each star, we extract stellar parameters (effective temperature $\rm T_{eff}$, $\rm\log{g}$), as well as [Fe/H] and [C/Fe]. Since Eu and Ba abundances are available for a very limited number of stars, further classification into CEMP-no, s and r/s is not straightforward. However, as shown by Yoon et al. (2018), classification into group I, II and III is still possible based on [C/Fe] and [Fe/H] alone, which then allows an approximate separation of CEMP-no stars from those that are enhanced in n-capture elements.
In total, our sample contains $3192$ stars with sub-solar metallicities, [Fe/H] <0<0, out of which $1745$ have a measured [C/Fe] abundance, with $541$ exceeding $\rm[C/Fe]\geq 0.7$ and hence can be classified as CEMP (Yoon et al. 2018). The references to individual datasets along with stellar parameters and abundances (original and $\rm 3D-corrected$ values) are provided as supplementary material via CDS.
@@ 2.2 Model atmospheres
In this work we rely on the M3DIS (Models in 3D@DISPATCH) code within the DISPATCH framework (Nordlund et al. 2018) for generating 3D RHD stellar atmospheres, which has been introduced and extensively tested by Eitner et al. (2024). In short: M3DIS solves the equations of hydrodynamics (HD) together with the radiative transfer (RT) equation to compute the radiative cooling, which in turn contributes to the energy balance of the atmosphere. The code relies on a HLLC Riemann solver for advancing the simulation in time (Fromang et al. 2006), and on short-characteristics radiative transfer in two $\rm\mu$ (polar) and four $\rm\phi$ (azimuthal) angles, plus upward and downward vertical rays, assembled in a Gauss-Radau quadrature. The main features of the models can be found in Table 1.
@@
The RT is solved in the multi-group approximation (Nordlund 1982; Ludwig et al. 1990), such that the monochromatic opacities and source functions are collected into eight bins based on wavelength and formation height within the respective atmosphere. The selection of opacity bin edges is done automatically using a generic k-means clustering approach. Monochromatic opacities are computed together with the internal energy and gas pressure, i.e., the equation of state (EoS), on a grid of $\rm 151\times 151$ $\rm\log\rho-\log T$ points using the updated version of MULTI3D (Leenaarts & Carlsson 2009), also within the DISPATCH framework. Scattering due to free electrons and Rayleigh scattering (H I, H2, He I) is treated as true absorption in the optically thick-part of the opacity table, and it is not excluded in the optically-thin regime. This is a common approach in 1D (Gerber et al. 2023) and 3D (Collet et al. 2011) calculations. Opacities and source functions are computed under the assumption of local-thermodynamic equilibrium (LTE) in the context of computing surface cooling. For more details see the description by Eitner et al. (2024), and Appendix B.
@@
Besides its computation speed, major advantages of M3DIS are the highly flexible and consistent opacities and EoS, which make the simulation of stars with non-solar chemical composition relatively straightforward. For investigating the impact of the carbon abundance on the atmospheric structure itself, we first compute scaled-solar $\rm[C/Fe]$ composition (relative to Magg et al. 2022) opacity and EoS tables at $\rm[Fe/H]=-6,-5,-4,-3,\ and\ -2$. Second, we calculate carbon enhanced opacity tables, such that at $\rm[Fe/H]\leq-3$ we additionally have C-enhanced atmospheres with the same, constant $\rm A(C)=6.56$, while at $\rm[Fe/H]=-2$ we keep the carbon enhancement fixed at $\rm[C/Fe]=1,\ A(C)=7.56$ (in agreement with the Galactic CEMP abundance trend, e.g. as reported by Yoon et al. 2016, Fig. 1). Because the computation of 3D RHD stellar atmosphere models is computationally expensive, previous studies mostly relied on a single, representative set of stellar parameters (e.g. Gallagher et al. 2017; Collet et al. 2018). To minimize uncertainties originating in the variety of stellar parameters in the data sample, we create solar-like MS models at $\rm T_{eff}=5750\ K,\ log(g)=4.5$, and cooler sub-giants at $\rm T_{eff}=5250\ K,\ log(g)=3.0$ for all chemical compositions. The chosen parameters represent a compromise between spanning the full range of stellar parameters and accurately reflecting the bulk of the distribution (see the discussion in appendix A for further information). To ensure a satisfactory resolution, the number of grid points has been set to $240\times 240\times 120$. The physical size of the simulation domain has been limited between the average Rosseland optical depth of $\rm\log\tau_{ross}=-6$ and +6.5+6.5, which sets the geometric domain of the box. This is done to ensure that the models are deep enough to recover the adiabaticity at the bottom of the atmosphere and the relevant line formation region of the CH G-band at the top. The resulting geometric domain ranges from approximately $2.5$ Mm for the MS 3D models (vertically) to 80 Mm for the sub-giant models. Horizontally, the models are twice as extended as vertically. M3DIS models are so-called “box-in-a-star” models, focusing only on the outermost layers of stars. Therefore, no additional assumptions about stellar parameters are necessary. For a typical mass of $\rm 0.8\,M_{\odot}$ (Ruchti et al. 2013), the MS model resembles a star with a radius of $\rm 0.83\,R_{\odot}$ and the sub-giant $\rm 4.7\,R_{\odot}$. A summary of the model parameters can be found in Tab. 2.
@@
For comparison, we use the classical 1D HE, scaled-solar, LTE model atmospheres from the MARCS grid (Gustafsson et al. 2008). To evaluate the effect of carbon enhancement also in 1D hydrostatic models, we furthermore obtained a carbon-enhanced MARCS model with stellar parameters $\rm T_{eff}=5750\ K,\ log(g)=4.5,\ [Fe/H]=-5,\ [C/Fe]=+3.0$ that we can include in the analysis (B. Plez, priv. conv.).
@@
Figure 1: Comparison of the mean temperature stratification between 1D scaled-solar MARCS (red solid lines), 3D scaled-solar (blue solid lines), and 3D CEMP (blue crosses) models against mean density (top) and optical depth at $\rm 500\ nm$ (bottom). Left: MS model. Right: sub-giant model. For the MS model in the left panels we furthermore include a CEMP MARCS model (provided by B. Plez, priv. conv., shown with red crosses)

### v2

# M3DIS - A grid of 3D radiation-hydrodynamics stellar atmosphere models for stellar surveys

@@ Abstract
Aims. Understanding the origin and evolution of carbon-enhanced metal-poor (CEMP) stars is key to tracing the early chemical enrichment of the Galaxy. In this work we investigate how physically realistic 3D radiation-hydrodynamic (RHD) carbon-enhanced model atmospheres affect the inferred carbon abundances in CEMP stars, and assess the implications for their classification and for Galactic chemical evolution (GCE). We pay particular attention to the systematic biases introduced by traditional 1D hydrostatic equilibrium (HE) models.
Methods. We used the M3DIS code to compute 3D RHD model atmospheres for main-sequence and sub-giant stars spanning a wide range of metallicities and carbon enhancements. Synthetic spectra of the CH G band were calculated using full 3D radiative transfer and compared to spectra from classical 1D HE MARCS models. We derived abundance corrections and applied them to a large literature sample of metal-poor stars from the SAGA database to quantify systematic effects on the carbon abundance distribution and CEMP classification.
Results. Our new 3D CEMP models predict significantly cooler upper atmospheric layers than 1D HE models, resulting in stronger CH absorption and lower inferred carbon abundances by up to 0.9-0.9 dex at the lowest metallicities. Carbon enhancement in the atmosphere itself increases molecular opacities and leads to radiative re-heating, which partly offsets the adiabatic cooling in 3D models and reduces the magnitude of $\rm 3D-1D$ abundance corrections. Applying these corrections lowers the CEMP fraction by up to $20\%$ below $\rm[Fe/H]=-3$, and furthermore alters the relative contribution of CEMP sub-classes. In particular, the fraction of stars classified as CEMP-no increases and that of CEMP-r/s stars decreases, owing to the downward revision of absolute carbon abundances. These changes bring the Galactic carbon abundance distribution into better agreement with GCE models that assume a contribution from faint supernovae of $20\%$. Physically realistic model atmospheres are thus essential for a reliable reconstruction of the early chemical enrichment history of the Galaxy.
@@ 1 Introduction
Carbon-enhanced metal-poor (CEMP) stars are of major interest in modern astrophysics. Chemical abundances measured in the spectra of these stars are relevant for understanding the properties of type II supernova (SN) events in the early Galaxy, nucleosynthesis, the formation and evolution of binary stars, the stellar initial mass function, the metallicity distribution function of the Galactic halo, and the production sites of neutron-capture elements (Beers & Christlieb 2005; Aoki et al. 2006; Hansen et al. 2018).
@@
Carbon-enhanced metal-poor stars are defined based on the amount of carbon (C) compared to iron (Fe) relative to the solar composition present in their atmospheres. Originally, the limit for CEMP classification was placed at $\rm[C/Fe]>1.0$ (Beers & Christlieb 2005), which was later slightly adjusted to the presently applied criterion of $\rm[C/Fe]>0.7$ (Aoki et al. 2007). It is furthermore common to subclassify CEMP stars based on the presence of elements synthesised in the slow or rapid neutron capture processes (s- and r-processes; Hirschi et al. 2006). The categories include stars with enhancements in one of the two channels (CEMP-s and CEMP-r), in both (CEMP-r/s), or in neither (CEMP-no), and each category is thought to probe a different formation environment (e.g. Masseron et al. 2010). CEMP-s stars are viewed as products of enrichment by a post-asymptotic giant branch companion in a binary system (McClure & Woodsworth 1990; Lucatello et al. 2005; Starkenburg et al. 2013; Hansen et al. 2016a), although some CEMP-s stars appear to be single, which might indicate that they formed from the ejecta of a massive, fast rotating source star with a late mixing event (Choplin 2017). In contrast, CEMP-r stars comprise a larger enhancement of r-process elements, such as Eu (Sneden et al. 2003), and are thought to form from matter that became chemically polluted by a core-collapse supernova (CCSN; electron-capture or magneto-rotating), a neutron star–neutron star merger, or a neutron star–black hole merger (Frebel & Norris 2015; Côté et al. 2017; Hansen et al. 2018; Kobayashi et al. 2020).
@@
The formation history of CEMP-r/s and CEMP-no stars is less clear. Regarding CEMP-r/s, there is evidence of an r-process contribution (Aoki et al. 2006) from a SN II. Alternative explanations include s-process enrichment by a thermally pulsating asymptotic giant branch companion (Masseron et al. 2006) and neutrino-driven winds from the accretion-induced collapse of an O-Ne-Mg white dwarf (Cohen et al. 2003; Jonsell et al. 2006; Kobayashi et al. 2020). CEMP-no stars are among the most metal-poor stars observed (Christlieb et al. 2002; Hansen et al. 2016b) and show no significant enrichment in either the r- or s-process (Spite et al. 2013). This can be interpreted in the framework of different scenarios, including nucleosynthesis in faint low-energy SNe (Umeda et al. 2005; Tominaga et al. 2007; Nomoto et al. 2013; Kobayashi et al. 2020), hypernovae (Heger & Woosley 2002; Limongi et al. 2003; Nomoto et al. 2013), nucleosynthesis in C-rich winds of rotating massive stars (Hirschi et al. 2006; Meynet et al. 2006), and early-asymptotic giant branch stars (Masseron et al. 2010, and references therein).
@@
The analysis of C abundances in the atmospheres of CEMP stars primarily rests upon spectroscopic diagnostic of the so-called G band (e.g. Placco et al. 2016; Placco & Beers 2016; François et al. 2018; Norris & Yong 2019). The G band represents the fundamental A-X transition of the CH molecule (Masseron et al. 2014; Popa et al. 2023) and is therefore highly sensitive to the abundance of C and to the temperature gradient in stellar atmospheres. The G band has been extensively studied using 1D local thermodynamic equilibrium (LTE) models (e.g. Lai et al. 2007; Roederer et al. 2014; Jacobson et al. 2015; Lee et al. 2013, see also Suda et al. 2008, 2017 and references therein), but recent analyses have also explored the formation of the G band in 3D radiation-hydrodynamics (RHD) model atmospheres of metal-poor dwarfs and red giants (e.g. Collet et al. 2007; Bonifacio et al. 2013; Gallagher et al. 2016, 2017; Collet et al. 2018).
@@
In this study we investigated the effect of carbon enhancement on the structure of stellar atmospheres in realistic 3D RHD simulations introduced in our previous work (Eitner et al. 2024). We further studied and quantified the impact of 3D convection on the diagnostics of C abundances from stellar spectra. We then analysed the results with respect to CEMP statistics in the Galaxy and discuss implications for the formation and evolution of CEMP stars.
@@
The paper is organised as follows. In Sect. 2 we briefly introduce the stellar observations used in this work, followed by a description of the M3DIS code and model parameters. In Sect. 3 we present our new 3D RHD model atmosphere calculations, including a comparison with 1D hydrostatic equilibrium (HE) MARCS models, the spectrum synthesis of the CH G band between 42974297\,Å and 43034303\,Å, and the analysis of CEMP abundances in 3D, along with a discussion of their implications for Galactic CEMP populations. Finally, in Sect. 4 we summarise our findings and discuss their implications for the study of CEMP stars and Galactic chemical evolution (GCE).
@@ 2 Methods
Table 1: M3DIS RHD code features.
@@
Radiative Transfer Short-characteristics; conservative on staggered mesh; 8 frequency bins; 4 $\rm\phi$, 2 $\rm\mu$ angles, plus disk-centre rays
@@ 2.1 Stellar observations
In this study we relied on stellar C abundances provided in the Stellar Abundances for Galactic Archaeology (SAGA) database [^1] http://sagadatabase.jp/ (Suda et al. 2008, 2011, 2017). SAGA is a comprehensive source of abundances and stellar parameters for stars in the Milky Way and its satellites, with a particular focus on metal-poor stars. The data are based on published, peer-reviewed sources and it has a particular focus on homogeneity of measurements, such that abundances are derived from 1D LTE synthetic spectra (Suda et al. 2008).
The focus of this study was main-sequence (MS) and sub-giant stars. Hence, we did not include stars with a surface gravity $\rm log(g)<2.7$. For each star, we extracted stellar parameters (the effective temperature and $\rm\log{g}$), as well as [Fe/H] and [C/Fe]. Since Eu and Ba abundances are available for a very limited number of stars, further classification into CEMP-no, s and r/s is not straightforward. However, as shown by Yoon et al. (2018), classification into group I, II and III is still possible based on [C/Fe] and [Fe/H] alone, which then allows an approximate separation of CEMP-no stars from those that are enhanced in n-capture elements.
In total, our sample contains $3192$ stars with sub-solar metallicities, [Fe/H] <0<0, out of which $1745$ have a measured [C/Fe] abundance, with $541$ exceeding $\rm[C/Fe]\geq 0.7$ and hence can be classified as CEMP (Yoon et al. 2018). The references to individual datasets along with stellar parameters and abundances (original and $\rm 3D-corrected$ values) are provided as supplementary material via the CDS.
@@ 2.2 Model atmospheres
In this work we relied on the M3DIS (Models in 3D@DISPATCH) code within the DISPATCH framework (Nordlund et al. 2018) for generating 3D RHD stellar atmospheres, which has been introduced and extensively tested by Eitner et al. (2024). In short: M3DIS solves the equations of hydrodynamics together with the radiative transfer (RT) equation to compute the radiative cooling, which in turn contributes to the energy balance of the atmosphere. The code relies on a HLLC Riemann solver for advancing the simulation in time (Fromang et al. 2006), and on short-characteristics RT in two $\rm\mu$ (polar) and four $\rm\phi$ (azimuthal) angles, plus upward and downward vertical rays, assembled in a Gauss-Radau quadrature. The main features of the models can be found in Table 1.
@@
The RT is solved in the multi-group approximation (Nordlund 1982; Ludwig et al. 1990), such that the monochromatic opacities and source functions are collected into eight bins based on wavelength and formation height within the respective atmosphere. The selection of opacity bin edges is done automatically using a generic k-means clustering approach. Monochromatic opacities are computed together with the internal energy and gas pressure, i.e. the equation of state (EoS), on a grid of $\rm 151\times 151$ $\rm\log\rho-\log T$ points using the updated version of MULTI3D (Leenaarts & Carlsson 2009), also within the DISPATCH framework. Scattering due to free electrons and Rayleigh scattering (H I, H2, and He I) is treated as true absorption in the optically thick-part of the opacity table, and it is not excluded in the optically thin regime. This is a common approach in 1D (Gerber et al. 2023) and 3D (Collet et al. 2011) calculations. Opacities and source functions are computed under the assumption of LTE in the context of computing surface cooling. For more details, see the description by Eitner et al. (2024) and Appendix B.
@@
In addition to its computation speed, major advantages of M3DIS are the highly flexible and consistent opacities and EoS, which make the simulation of stars with non-solar chemical composition relatively straightforward. To investigate the impact of the carbon abundance on the atmospheric structure itself, we first computed the scaled-solar $\rm[C/Fe]$ composition (relative to Magg et al. 2022) opacity and EoS tables at $\rm[Fe/H]=-6,-5,-4,-3,\ and\ -2$. Second, we calculated carbon-enhanced opacity tables, such that at $\rm[Fe/H]\leq-3$ we additionally have C-enhanced atmospheres with the same, constant $\rm A(C)=6.56$, while at $\rm[Fe/H]=-2$ we keep the carbon enhancement fixed at $\rm[C/Fe]=1,\ A(C)=7.56$ (in agreement with the Galactic CEMP abundance trend, e.g. as reported by Yoon et al. 2016, Fig. 1). Because the computation of 3D RHD stellar atmosphere models is computationally expensive, previous studies mostly relied on a single, representative set of stellar parameters (e.g. Gallagher et al. 2017; Collet et al. 2018). To minimise uncertainties originating in the variety of stellar parameters in the data sample, we created solar-like MS models at $\rm T_{eff}=5750\ K,\ log(g)=4.5$, and cooler sub-giants at $\rm T_{eff}=5250\ K,\ log(g)=3.0$ for all chemical compositions. The chosen parameters represent a compromise between spanning the full range of stellar parameters and accurately reflecting the bulk of the distribution (see the discussion in Appendix A for further information). To ensure a satisfactory resolution, the number of grid points has been set to $240\times 240\times 120$. The physical size of the simulation domain has been limited between the average Rosseland optical depth of $\rm\log\tau_{ross}=-6$ and +6.5+6.5, which sets the geometric domain of the box. This is done to ensure that the models are deep enough to recover the adiabaticity at the bottom of the atmosphere and the relevant line formation region of the CH G band at the top. The resulting geometric domain ranges from approximately $2.5$ Mm for the MS 3D models (vertically) to 80 Mm for the sub-giant models. Horizontally, the models are twice as extended as vertically. M3DIS models are so-called box-in-a-star models, focusing only on the outermost layers of stars. Therefore, no additional assumptions about stellar parameters are necessary. For a typical mass of $\rm 0.8\,M_{\odot}$ (Ruchti et al. 2013), the MS model resembles a star with a radius of $\rm 0.83\,R_{\odot}$ and the sub-giant $\rm 4.7\,R_{\odot}$. A summary of the model parameters can be found in Table 2.
@@
For comparison, we used the classical 1D HE, scaled-solar, LTE model atmospheres from the MARCS grid (Gustafsson et al. 2008). To evaluate the effect of carbon enhancement also in 1D hydrostatic models, we furthermore obtained a carbon-enhanced MARCS model with stellar parameters $\rm T_{eff}=5750\ K,\ log(g)=4.5,\ [Fe/H]=-5,and\ [C/Fe]=+3.0$ to use in the analysis (B. Plez, priv. conv.).
@@
Figure 1: Comparison of the mean temperature stratification between 1D scaled-solar MARCS (solid red lines), 3D scaled-solar (solid blue lines), and 3D CEMP (blue crosses) models against mean density (top) and optical depth at $\rm 500\ nm$ (bottom). Left: MS model. We include a CEMP MARCS model (provided by B. Plez, priv. conv., shown with red crosses). Right: Sub-giant model.